Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 191     Entry time: 07 May 2023, 06:52
Author: Giuseppe Mercurio 
Group: SCS 
Subject: shift summary 

achievements:

  • measured delay scans on 16 nm samples at different magnetic fields (4 runs) before the problem below appeared

problems:

  • scroll pump of the FFT degraded, rough vacuum got worse, this led to pump in error state and interlock valves. The scroll pump was replaced, vacuum re-established, but we could not see any signal from the samples, only diffuse scattering. We thought this was due to ice frozen on the samples when the pressure increased. We warmed up the cryostat, verified that the conditions of the samples are fine again, and finally started cooling down again.
ELOG V3.1.4-7c3fd00